OpenStudy (anonymous):

PLEASE HELP! Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2− 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (stamp):

I am inclined to say b^2-4ac, also known as the discriminant.

OpenStudy (anonymous):

can you show me how you got that answer?

OpenStudy (stamp):

\[\frac{-b+/-\sqrt(b^2-4ac)}{2a}\]

OpenStudy (stamp):

if the discriminant is > 0, the polynomial has two real roots, if the discriminant is = 0, the polynomial has one real double root, and if the discriminant is < 0, the polynomial has no real roots.

OpenStudy (stamp):

\[2x^2 + 7x + 3 = 0 \] So the discriminant is 49 - 24 > 0, meaning the polynomial has two real roots.